# Sample Complexity of Diffusion Model Training Without Empirical Risk Minimizer Access

### v2

@@ 4 Conclusion and Future Work
In this work, we investigate the sample complexity of training diffusion models via score estimation using neural networks. We derive a sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-6})$, which, to our knowledge, is the first such result that does not assume access to an empirical risk minimizer of the score estimation loss. Notably, our bound does not depend exponentially on the number of neural network parameters. For comparison, the best-known existing result achieves a bound of $\widetilde{\mathcal{O}}(\epsilon^{-5})$, but it crucially assumes access to an ERM. All prior results establishing sample complexity bounds for diffusion models have made this assumption. Our contribution is the first to establish a sample complexity bound for diffusion models under the more realistic setting where exact access to empirical risk minimizer of the score estimation loss is not available.
While our analysis focuses on continuous distributions, extending these guarantees to discrete settings remains an important direction for future work. A promising direction would be to extend our results to conditional diffusion models such as classifier free and classifier guided diffusion.
Appendix A Appendix
This appendix contains additional technical material, including full proofs of our main results, detailed explanations of key lemmas, algorithmic descriptions, and further discussion to support and extend the main text.
Appendix B Comparison with Prior Works
In this section, we provide a detailed comparison of our results with prior work. Specifically, we analyze the sample complexity bounds presented in [15], and show how combining their results with those of [7] leads to an alternative bound.
B.1 Sample Complexity of [15]
We begin by examining the sample complexity claim of $\mathcal{O}(1/\epsilon^{3})$ reported in [15]. A closer analysis reveals that the actual sample complexity is $\widetilde{\mathcal{O}}(1/\epsilon^{5})$, once the error over all the discretization steps is properly accounted for and a union bound is applied.
The main result regarding the sample complexity of estimating the score function as given in Theorem C.2 of [15] is as follows. We re-iterate this Theorem here.
Theorem C.2 [15]. Let $q$ be a distribution of $\mathbb{R}^{d}$ with second moment $m_{2}^{2}$. Let $\phi_{\theta}(\cdot)$ be the fully connected neural network with ReLU activations parameterized by $\theta$, with $P$ total parameters and depth $D$. Let $\Theta>1$. For any $\gamma>0$, there exist $K=\widetilde{\mathcal{O}}\left(\frac{d}{\epsilon^{2}+\delta^{2}}\log^{2}\frac{m_{2}+1/m_{2}}{\gamma}\right)$ discretization times $0=t_{0}<\cdots<t_{K}<T$ such that if for each $t_{k}$, there exists some score function $\hat{s}_{\theta}$ with $\|\theta^{*}\|_{\infty}\leq\Theta$ such that
@@ B.1 Sample Complexity of [15]
Note that Lemma B.6 of [15] states that
$\mathrm{TV}(p,\hat{p})\leq{\delta}+\mathcal{O}\left(\frac{1}{\sqrt{N}}\right)+{\epsilon}\cdot{\sqrt{T}}+\mathcal{O}(\exp^{-T})$ (34)
@@
In order to achieve $\mathrm{TV}(p,\hat{p})\leq\epsilon$ we have to set $\delta=\epsilon$. This would imply $N=\mathcal{O}(\epsilon^{-2})$ and $T=\mathcal{O}\left(\frac{1}{\epsilon}\right)$. Putiing this value of $N$ in Equation (33) we obtain that for
@@
In contrast, our analysis reduces the overall complexity by a factor of $\widetilde{\mathcal{O}}(1/\epsilon)$, yielding the tightest known bounds for neural score estimation in diffusion models, i.e., $\widetilde{\mathcal{O}}(1/\epsilon^{4})$. Further, unlike [15], our analysis avoids using the $1-\delta$-quantile bound on the score norm and instead directly bounds the global $L^{2}$ score estimation error thus avoids applying a union bound across time steps, and finally achieve tighter sample complexity guarantees.
B.2 A note on the sample complexity result in [7]
The authors of [15] cite the intermediate results of [7] which obtained a sample complexity result of $\tilde{O}(d^{\frac{5}{2}}\cdot W^{D/2}\cdot\epsilon^{-2})$ for a single discretization point. In Section 4 of [15] it is explained that simply obtaining an upper bound on the L2 error at one discretization point is not sufficient to obtain a TV bound between the learned and true probability distributions with probability. This is formally stated in Lemmas 4.1 and 4.2 where it is proved that even when the true distribution is a simple Gaussian, there is a lower bound on the L2 error between the true and estimated score function. Thus the sample complexity of $\tilde{O}(d^{\frac{5}{2}}\cdot W^{D/2}\cdot\epsilon^{-2})$ proved in [7] does not ensure convergence in TV between the true and estimated probability distributions. Additionally, the sample complexity result in [7] is exponential in the depth of of the score estimation neural network, something which is not the case in the sample complexity result we have achieved in this work.
Appendix C Score Estimation Algorithm
@@
3:  for $k=1$ to $n$ do
4:     Sample $x_{i}\sim\mathcal{D}$, $t_{i}\sim\mathrm{Uniform}([1,T])$, $\epsilon_{i}\sim\mathcal{N}(0,I)$ for $i=1,\dots,B$
5:     $x_{t_{i}}=\sqrt{\bar{\alpha}_{t_{i}}}x_{i}+\sqrt{1-\bar{\alpha}_{t_{i}}}\epsilon_{i}$
6:     Compute loss: $\hat{L}(\theta)=\|\epsilon_{i}-\epsilon_{\theta}(x_{t_{i}},t_{i})\|^{2}$
7:     Update $\theta\leftarrow\theta-\eta_{k}\cdot\nabla_{\theta}\hat{L}(\theta)$
8:  end for Sampling
9:  Sample $x_{T}\sim\mathcal{N}(0,I)$
10:  for $t=T$ down to $t_{\text{stop}}+1$ do
11:     $z\sim\mathcal{N}(0,I)$ if $t>1$ else $z=0$
12:     $\hat{\epsilon}=\epsilon_{\theta}(x_{t},t)$
13:     $\tilde{\mu}_{t}=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\epsilon}\right)$
14:     $x_{t-1}=\tilde{\mu}_{t}+\sqrt{\beta_{t}}\cdot z$
15:  end for
16:  Return $x_{t_{\text{stop}}}$
Appendix D Proofs of Intermediate Lemmas
@@
In this section, we present the proofs of intermediate lemmas used to bound the approximation error, statistical error, and optimization error in our analysis.
D.1 Bounding the Approximation Error
Recall Lemma 1 (Approximation Error.) Let $W$ and $D$ denote the width and depth of the neural network architecture, respectively, and let $d$ represent the input data dimension. Then, under Assumption 3 for all $k\in[{0},K]$ we have
$\mathcal{E}^{\mathrm{approx}}_{t}\leq\mathcal{O}\left(\left(WD\right)^{-\frac{2}{d}}\right).$ (37)
Proof.
Recall that the approximation error $\mathcal{E}^{\mathrm{approx}}_{t}$ captures the inherent expressiveness limitations of the function class parameterized by the neural network. Even with infinite data and perfect optimization, the network may not exactly represent the true score function $s^{*}(x)=\nabla\log p_{t}(x)$. This error is defined as:
$\mathcal{E}^{\text{approx}}_{t}=\min_{\theta\in\Theta}\mathbb{E}_{x\sim p_{t}}\left\|s_{\theta}(x)-\nabla\log p_{t}(x)\right\|^{2},$ (38)
where $\Theta$ denotes the parameter space of the neural network class.
To apply known approximation results for bounded target functions, we first analyze the probability that the score function $\nabla\log p_{t}(x)$ is bounded in norm. For a fixed threshold $\kappa>0$, consider
$\mathbb{P}\left(\left\|\nabla\log p_{t}(x)\right\|\geq\kappa\right).$ (39)
Using the law of total expectation, we have
$\mathbb{E}_{p}\left[\mathbb{P}\left(\left\|\nabla\log p_{t}(x)\right\|\geq\kappa\,\middle|\,x_{0}\right)\right].$ (40)
Now, conditioned on the data $p$, the forward process implies that $x_{t}$ follows a Gaussian distribution: $x_{t}\sim\mathcal{N}(\alpha_{t},\sigma^{2}_{t}{\cdot}I)$. Therefore, for fixed $x_{0}$
$\mathbb{P}\left(\left\|\nabla\log p_{t}(x)\right\|\leq\kappa\,\middle|\,x_{0}\right)=\mathbb{P}\left(\left\|\frac{x-\alpha_{t}}{\sigma_{t}}\right\|\geq\kappa\,\middle|\,x_{0}\right).$ (41)
The quantity $\frac{x-\alpha_{t}}{\sqrt{1-\alpha_{t}}}$ is Gaussian with mean 00 under this condition thus the expectation with respect to $p$ can be ignored. So we have
$\mathbb{P}\left(\left\|\nabla\log p_{t}(x)\right\|\geq\kappa\right)\leq Z_{\kappa,t}$ (42)
This implies
$\mathbb{P}\left(\left\|\nabla\log p_{t}(x)\right\|\leq\kappa\right)\leq Z_{\kappa}$ (43)
where the constant $Z_{\kappa,t}$ does not depend on $x_{{}_{0}}$. $Z_{\kappa}=\sup_{t\in[0,T]}Z_{\kappa,t}$ which ensures that this bound holds for all $t\in[0,T]$ denotes the tail probability of the standard normal distribution.
We can now fix $\kappa$ high enough such that that $Z_{\kappa}\leq\frac{\delta}{2}$, ensuring that with probability at least $1-\delta$, the score function is bounded in norm by $\kappa$. Now note that the expectation $E_{x\sim p_{t}}||s_{\theta}-\nabla{\log}p_{t}(x)||^{2}$ is bounded. Therefore, there exists an unbounded set $K$ such that $P(x\in K)\leq\frac{\delta}{2}$ and the complement of $K$ is a bounded set. Thus we get
$P(A\cup K)\leq P(A)+P(K)\leq{\delta}$ (44)
$P(\bar{A}\cap\bar{K})\geq 1-\delta$ (45)
Thus we have with probability at least $1-\frac{\delta}{2}$ that the function ${\nabla}{\log}p_{t}(x)$ is bounded and belongs to a bounded set. we invoke Theorem 2.5 from [36], which states that there exists a score function $s_{\tilde{\theta}}$ which is a neural network with ReLU activation such that.
$||{\nabla}{\log}p_{t}(x)-s_{\tilde{\theta}}||_{\infty}\leq(W.D)^{-\frac{2}{d}}$ (46)
Now Theorem 6 of [41] states that for any neural network with ReLU activation $s_{\tilde{\theta}}$ there exists a neural network with a smooth activation such that.
$||s_{\tilde{\theta}}-s_{{\theta}}||_{\infty}\leq\epsilon$ (47)
for any $\epsilon>0$. Now setting $\epsilon=(W.D)^{-\frac{2}{d}}$ we get that there exists a neural network with a smooth activation function such that
$||{\nabla}{\log}p_{t}(x)-s_{\tilde{\theta}}||_{\infty}$ $\leq||{\nabla}{\log}p_{t}(x)-s_{\tilde{\theta}}||_{\infty}+||s_{\tilde{\theta}}-s_{{\theta}}||_{\infty}$ (48)
$\leq(W.D)^{-\frac{2}{d}}$ (49)
Since the bound holds for $||.||_{\infty}$ we can take expectation with respect to ${p_{t}}$ on both sides of Equation (49) to get the required result. ∎
D.2 Bounding the Statistical Error
Next, we bound the statistical error $\mathcal{E}_{\text{stat}}$. This error is due to estimating $s^{*}(x)$ from a finite number of samples rather than the full data distribution.
Recall Lemma 2 (Statistical Error.) Let $n$ denote the number of samples used to estimate the score function at each diffusion step. Then, for all $k\in[0,K]$, with probability at least $1-\delta$, we have
$\mathcal{E}^{\mathrm{stat}}_{k}\leq\mathcal{O}\left(d.\sqrt{\frac{{{\log}\left(\frac{2}{\delta}\right)}}{{n}}}\right)$ (51)
@@ D.2 Bounding the Statistical Error
$\mathcal{L}_{k}(\theta)=\mathbb{E}_{x\sim p_{t_{K}}}\left\|s_{\theta}(x,t)-\nabla\log p_{t}(x)\right\|^{2},$ (52)
where $s_{\theta}$ denotes the score function estimated by a neural network parameterized by $\theta$, and $x$ denotes samples at time $t$ used in Algorithm 1.
The corresponding empirical loss is defined as:
^$\widehat{\mathcal{L}}_{k}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\left\|s_{\theta}(x_{i},t)-\nabla\log p_{t}(x_{i})\right\|^{2}.$ (53)
Let $\theta^{a}_{k}$ and $\theta^{b}_{k}$ be the minimizers of $\mathcal{L}_{k}(\theta)$ and ^$\widehat{\mathcal{L}}_{t}(\theta)$, respectively, corresponding to score functions $s^{a}_{t_{k}}$ and $s^{b}_{t_{k}}$. By the definitions of minimizers, we can write
@@
Note that right hand side of Equation (54) is less then the left hand side since we have added the quantity ^$\widehat{\mathcal{L}}_{t}(\theta^{a}_{t})-\widehat{\mathcal{L}}_{t}(\theta^{b}_{t})$ which is strictly positive since $\theta^{b}_{k}$ is the minimizer of the function ^$\widehat{\mathcal{L}}_{t}(\theta)$ by definition. We then take the absolute value on both sides of Equation (55) to get
$|\mathcal{L}_{k}(\theta^{b}_{k})-\mathcal{L}_{k}(\theta^{a}_{k})|$ $\leq\underbrace{\left|\mathcal{L}_{k}(\theta^{b}_{k})-\widehat{\mathcal{L}}_{t}(\theta^{b}_{k})\right|}_{\text{(I)}}+\underbrace{\left|\mathcal{L}_{k}(\theta^{a}_{k})-\widehat{\mathcal{L}}_{t}(\theta^{a}_{k})\right|}_{\text{(II)}}.$ (56)
Note that this preserves the direction of the inequality since the left hand side of Equation (55) is strictly positive since $\theta^{a}_{k}$ is the minimizer of $\mathcal{L}_{k}(\theta)$ by definition.
@@
We now bound terms (I) and (II) using standard generalization results. From Theorem 26.5 of [35], if the loss function is uniformly bounded over the parameter space $\Theta^{\prime\prime}=\{\theta^{a}_{t},\theta^{b}_{t}\}$, then with probability at least $1-\delta$, we have
@@
$\mathbb{E}\left[\left|\mathcal{L}_{k}(\theta)-\widehat{\mathcal{L}}_{t}(\theta)\right|\right]\leq\widehat{R}(\theta)+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right),~{}~{}\forall~{}\theta\in\Theta^{\prime\prime}$ (57)
@@
where $\widehat{R}(\theta)$ denotes the empirical Rademacher complexity of the function class restricted to $\Theta^{\prime\prime}$.
To ensure the boundedness of the loss, we define the event $A$ as
$A:=\left\{x:\left\|\nabla\log p_{t}(x)\right\|\leq\kappa\right\}.$ (58)
As shown in Lemma 1, $\mathbb{P}(A)\geq 1-\delta$ for an appropriately large constant $\kappa$, based on the Gaussian nature of $x_{t}$. Let event $B$ denote the failure of the generalization bound
$B:=\left\{\mathbb{E}\left[\left|\mathcal{L}_{k}(\theta)-\widehat{\mathcal{L}}_{t}(\theta)\right|\right]>\widehat{R}(\Theta^{\prime\prime})+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right)\right\}.$ (59)
From the above theorem, we know $\mathbb{P}(B)\leq\delta$ under the boundedness condition. Therefore, by the union bound, we have
$\mathbb{P}(A^{c}\cup B)$ $\leq\mathbb{P}(A^{c})+\mathbb{P}(B)\leq 2\delta,$ (60)
$\Rightarrow\mathbb{P}(A\cap B^{c})$ $\geq 1-2\delta.$ (61)
On this event $(A_{\kappa}\cap B^{c})$, both the score function is bounded and the generalization bound holds. Hence, with probability at least $1-2\delta$, we have:
@@
$\mathbb{E}\left[\left|\mathcal{L}_{k}(\theta)-\widehat{\mathcal{L}}_{t}(\theta)\right|\right]\leq\widehat{R}(\theta)+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right),~{}~{}\quad\forall~{}\theta\in\Theta^{\prime\prime}.$ (62)
Since $\Theta^{\prime\prime}$ is a finite class (just two functions). Note that Massart’s Lemma states that for a finite function class $\Theta^{{}^{\prime\prime}}$ we have
$\mathbb{E}_{\sigma}\max_{\theta\in\Theta^{{}^{\prime\prime}}}\sum_{i=1}^{n}f(\theta){\sigma_{i}}\leq||f(\theta)||_{2}\leq d.||\Theta||$ (63)
We get the second inequality by denoting $d$ as the dimension of $f(\theta)$ and $||\Theta||$ be a quantity such that $||\Theta||=\sup_{\theta\in\Theta^{{}^{\prime\prime}}}||\theta||_{\infty}$. Also, $\sigma=\{{\sigma_{1}},{\sigma_{2}},{\cdots},\sigma_{n}\}$
where $\sigma_{i}$ are i.i.d random variables such that $P(\sigma_{i}=1)=P(\sigma_{i}=-1)=\frac{1}{2}$.
Now, the definition of $\hat{R}(\theta)$ is $\hat{R}(\theta)=\frac{1}{m}\mathbb{E}_{\sigma}\max_{\theta\in\Theta^{{}^{\prime\prime}}}\sum_{i=1}^{n}f(\theta){\sigma_{i}}$. Thus by applying Massart’s lemma to $\hat{R}(\theta)$ in Equation (62) we get with probability at least $1-2.\delta$
$\mathbb{E}\left[\left|\mathcal{L}_{k}(\theta)-\widehat{\mathcal{L}}_{t}(\theta)\right|\right]\leq\mathcal{O}\left(\frac{d.||\Theta||}{n}\right)+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right),~{}~{}\quad\forall~{}\theta\in\Theta^{\prime\prime}.$ (64)
Here $||\Theta||$ is a value such that $||\Theta||=\sup_{\theta\in\Theta^{{}^{\prime\prime}}}||\theta||_{\infty}$. Which implies that $\Theta$ is the norm of the maximum parameter value that can be attained in the finite function class $\Theta^{{}^{\prime\prime}}$. which yields that with probability at least $1-2.\delta$ we have

### v5

@@ 4 Conclusion and Future Work
In this work, we investigate the sample complexity of training diffusion models via score estimation using neural networks. We derive a sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-4})$, which, to our knowledge, is the first such result that does not assume access to an empirical risk minimizer of the score estimation loss. Notably, our bound does not depend exponentially on the number of neural network parameters. For comparison, the best-known existing result achieves a bound of $\widetilde{\mathcal{O}}(\epsilon^{-5})$, but it crucially assumes access to an ERM. All prior results establishing sample complexity bounds for diffusion models have made this assumption. Our contribution is the first to establish a sample complexity bound for diffusion models under the more realistic setting where exact access to empirical risk minimizer of the score estimation loss is not available.
While our analysis focuses on unconditional distributions, extending these guarantees to conditional settings remains an important direction for future work.
Appendix A Comparison with Prior Works
In this section, we provide a detailed comparison of our results with prior work. Specifically, we analyze the sample complexity bounds presented in Gupta et al. (2024), and show how combining their results with those of Block et al. (2020) leads to an alternative bound.
A.1 Sample Complexity of Gupta et al. (2024)
We begin by examining the sample complexity claim of $\mathcal{O}(1/\epsilon^{3})$ reported in Gupta et al. (2024). A closer analysis reveals that the actual sample complexity is $\widetilde{\mathcal{O}}(1/\epsilon^{5})$, once the error over all the discretization steps is properly accounted for and a union bound is applied.
The main result regarding the sample complexity of estimating the score function as given in Theorem C.2 of Gupta et al. (2024) is as follows. We re-iterate this Theorem here.
Theorem C.2 Gupta et al. (2024). Let $q$ be a distribution of $\mathbb{R}^{d}$ with second moment $m_{2}^{2}$. Let $\phi_{\theta}(\cdot)$ be the fully connected neural network with ReLU activations parameterized by $\theta$, with $P$ total parameters and depth $D$. Let $\Theta>1$. For any $\gamma>0$, there exist $K=\widetilde{\mathcal{O}}\left(\frac{d}{\epsilon^{2}+\delta^{2}}\log^{2}\frac{m_{2}+1/m_{2}}{\gamma}\right)$ discretization times $0=t_{0}<\cdots<t_{K}<T$ such that if for each $t_{k}$, there exists some score function $\hat{s}_{\theta}$ with $\|\theta^{*}\|_{\infty}\leq\Theta$ such that
@@ Appendix A Comparison with Prior Works
Note that Lemma B.6 of Gupta et al. (2024) states that
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})\leq{\delta}+\mathcal{O}\left(\frac{1}{\sqrt{N}}\right)+{\epsilon}\cdot{\sqrt{T}}+\mathcal{O}(\exp^{-T})$ (35)
@@
In order to achieve $\mathrm{TV}(p,\hat{p})\leq\epsilon$ we have to set $\delta=\epsilon$. This would imply $N=\mathcal{O}(\epsilon^{-2})$ and $T=\mathcal{O}\left(\frac{1}{\epsilon}\right)$. Putiing this value of $N$ in Equation equation 34 we obtain that for
@@
In contrast, our analysis reduces the overall complexity by a factor of $\widetilde{\mathcal{O}}(1/\epsilon)$, yielding the tightest known bounds for neural score estimation in diffusion models, i.e., $\widetilde{\mathcal{O}}(1/\epsilon^{4})$. Further, unlike Gupta et al. (2024), our analysis avoids using the $1-\delta$-quantile bound on the score norm and instead directly bounds the global $L^{2}$ score estimation error thus avoids applying a union bound across time steps, and finally achieve tighter sample complexity guarantees.
Appendix B Score Estimation Algorithm
@@
3:for $k=1$ to $K$ do
4:  for $i=1$ to $n$ do
5:   Sample $x_{j}\sim\mathcal{D}$, $\epsilon_{k}\sim\mathcal{N}(0,I)$ for $i=1,\dots,n$
6:   $x_{t_{i}}=e^{-{t_{k}}}x_{i}+\sqrt{1-e^{-2t_{k}}}\epsilon_{i}$
7:   Compute loss: $\hat{L}(\theta)=\|\epsilon_{i}-\epsilon_{\theta}(x_{t_{i}},t_{i})\|^{2}$
8:   Update $\theta\leftarrow\theta-\eta_{i}\cdot\nabla_{\theta}\hat{L}(\theta)$
9:  end for
10:end forSampling
11: Sample $x_{T}\sim\mathcal{N}(0,I)$
12:for $t=T$ down to $t_{\text{stop}}+1$ do
13:  $z\sim\mathcal{N}(0,I)$ if $t>1$ else $z=0$
14:  $\hat{\epsilon}=\epsilon_{\theta}(x_{t},t)$
[…]
16:  $x_{t-1}=\tilde{\mu}_{t}+\sqrt{\beta_{t}}\cdot z$
17:end for
18:Return $x_{t_{\text{stop}}}$
Appendix C Proofs of Intermediate Lemmas
@@
In this section, we present the proofs of intermediate lemmas used to bound the statistical error and optimization error in our analysis.
C.1 Bounding the Statistical Error
@@
$\mathcal{L}_{k}(\theta)=\mathbb{E}_{x\sim p_{t_{k}}}\left\|s_{\theta}(x,t_{k})-\nabla\log p_{t_{k}}(x)\right\|^{2},$ (38)
where $s_{\theta}$ denotes the score function estimated by a neural network parameterized by $\theta$, and $x$ denotes samples at time $t$ used in Algorithm 1. The corresponding empirical loss is defined as:
$\widehat{\mathcal{L}}_{k}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\left\|s_{\theta}(x_{i},t_{k})-\nabla\log p_{t_{k}}(x_{i})\right\|^{2}.$ (39)
Let $\theta^{a}_{k}$ and $\theta^{b}_{k}$ be the minimizers of $\mathcal{L}_{k}(\theta)$ and $\widehat{\mathcal{L}}_{t}(\theta)$, respectively, corresponding to score functions $s^{a}_{t_{k}}$ and $s^{b}_{t_{k}}$. By the definitions of minimizers, we can write
@@
We now bound terms (I) and (II) using generalization results. From Lemma 5 (Theorem 26.5 of Shalev-Shwartz & Ben-David (2014)), if the loss function $\widehat{\mathcal{L}}(\theta)$ is uniformly bounded over the parameter space $\Theta^{\prime\prime}=\{\theta_{k}^{a},\theta_{k}^{b}\}$, then with probability at least $1-\delta$, we have
@@
$\left|\mathcal{L}_{k}(\theta)-\widehat{\mathcal{L}}_{t}(\theta)\right|\leq\widehat{R}(\Theta^{{}^{\prime\prime}})+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right),~~\forall~\theta\in\Theta^{\prime\prime}$ (42)
@@
where $\widehat{R}(\Theta^{{}^{\prime\prime}})$ denotes the empirical Rademacher complexity of the function class restricted to $\Theta^{\prime\prime}$. Now since $x$ is not bounded, this result does not hold. We then define the following two functions
$\mathcal{L^{\prime}}_{k}(\theta)=\mathbb{E}_{x\sim\mu_{t_{k}}}\left\|v_{\theta}(x,t_{k})-v_{t_{k}}(x)\right\|^{2},$ (43)
and
$\widehat{\mathcal{L^{\prime}}}_{k}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\left\|v_{\theta}(x_{i},t_{k})-v_{t_{k}}(x_{i})\right\|^{2}.$ (44)
where we define the functions
[…]
and
$\left(v_{\theta}(x,t)\right)_{j}=\begin{cases}\left(s_{\theta}(x,t)\right)_{j}&\text{if }|\frac{x-e^{-t}x_{0}}{\sigma^{2}_{t}}|_{j}\leq\kappa\\ 0&\text{if }|\frac{x-e^{-t}x_{0}}{\sigma^{2}_{t}}|_{j}\geq\kappa\end{cases}$ (46)
Here $\left(v_{t}(x)\right)_{j}$$\left(\nabla\log p_{t}(x)\right)_{j}$, $\left(v_{\theta}(x,t)\right)_{j}$ and $\left(s_{\theta}(x,t)\right)_{j}$ denote the $j^{th}$ co-ordinate of $v_{t}(x)$, $\left(\nabla\log p_{t}(x)\right)$, $v_{\theta}(x,t)$ and $s_{\theta}(x,t)$ respectively. Further, $|\frac{x-e^{-t}x_{0}}{\sigma^{2}_{t}}|_{j}$ denotes the $j^{th}$ co-ordinate of the $i^{th}$ sample of the score function in $\hat{\mathcal{L}_{k}}(\theta)$ which is given by ${\log}p_{t}(x)=|\frac{x-e^{-t}x_{0}}{\sigma^{2}_{t}}|$.
Note that the functions $v_{t}(x)$ and $v_{\theta}(x,t)$ are uniformly bounded. Thus using Theorem 26.5 of Shalev-Shwartz & Ben-David (2014) we have with probability at least $1-\delta$,
@@
$\left|\mathcal{L^{\prime}}_{k}(\theta)-\widehat{\mathcal{L^{\prime}}}_{k}(\theta)\right|\leq\widehat{R}(\Theta)+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right),~~\quad\forall~\theta\in\Theta^{\prime\prime}.$ (47)
Since $\Theta^{\prime\prime}=\{\theta_{a},\theta_{b}\}$ is a finite class (just two functions). We can apply Lemma E.5 to bound the empirical Rademacher complexity $\widehat{R}(\theta)$ in terms of the Rademacher complexity $R(\theta)$ of the function class $\Theta^{\prime\prime}$. Since $\widehat{R}(\theta)=\frac{1}{m}\mathbb{E}_{\sigma}\left[\max_{\theta\in\Theta^{{}^{\prime\prime}}}\sum_{i=1}^{n}f(\theta){\sigma_{i}}\right]$, applying Lemma E.5, we have with probability at least $1-2\delta$